OpenStudy (anonymous):

2n C 0 + 2n C 1 + 2n C 2 +....................2n C n = 2^(2n-1)..is it true. if yes or no, how??

OpenStudy (anonymous):

What is C? Some constant? It's missing from the RHS..

OpenStudy (anonymous):

combination sign..

OpenStudy (anonymous):

n choose k?

OpenStudy (anonymous):

n C r = {n!/(n-r)!r!} now, can u understand my question?

OpenStudy (anonymous):

yes, it's n choose k

OpenStudy (anonymous):

i am not getting you...

OpenStudy (anonymous):

It's the same thing, n choose k is a combination e.g. binomial coefficient

OpenStudy (anonymous):

it's just another way of saying it :)

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

It is false. It should be 2^n I think.

OpenStudy (anonymous):

Take n=1 then 4 c 0 + 4 c 1 = 1 + 4 = 5 but 2^(2-1) is not 5.

OpenStudy (anonymous):

No it should not be 2^n

OpenStudy (anonymous):

You are correct it is not 2^n. I was thinking sum k=0 to n of n c k but that expression is sum k=0 to n of 2n c k

OpenStudy (anonymous):

2^(2n-1).

OpenStudy (anonymous):

I was speculating what the correct formula was...I said 2^n since sum k=0 to n of n c k = 2^n is a common identity but I forgot your expression was 2n c k.

OpenStudy (anonymous):

Just ignore that post :) The identity is false. Just take n=1 for a counter example.